Find the mean, μ, for the binomial distribution which has the stated values of n and p. Round the answer to the nearest tenth. 19) n = 38; p = 3/5 (SHOW WORK) FINAL ANSWER:

Assume that a procedure yields a binomial distribution with a trial repeated n times. Use the binomial probability formula to find the probability of x successes given the probability p of success on a single trial. Round to three decimal places. 20) n = 5, x = 2, p = 0.70 (SHOW WORK) FINAL ANSWER:
Answer #1

1)

Mean = n p = 38 * 3 / 5 = 22.8

2)

X ~ bin ( n , p)

Where n = 5 , p = 0.7

Binomial probability distribution is

P(X) = nCx * px * ( 1 - p)n-x

P(X = 2) = 5C2 * 0.702 * ( 1 - 0.70)3

= 0.132